OpenStudy (anonymous):

x = 2 and y = 5 is a solution of: A. x + 2y = 12 B. x + y = 12 C. 2x + y = 10 D. x - 2y = 10

hartnn (hartnn):

just plug in x=2 and y=5 in each options

hartnn (hartnn):

example B : x+y = 2+5 = 7 but right side is 12 so B is incorrect

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

2+2*5=10 is it A?

hartnn (hartnn):

yes!

hartnn (hartnn):

2+2*5 = 12

OpenStudy (anonymous):

thanx champ

hartnn (hartnn):

welcome ^_^